# Nonreciprocal tripartite entanglement and asymmetric Einstein-Podolsky-Rosen steering via directional quantum squeezing

### v1

@@ II Theoretical model and system dynamics
where $(r,s,t)$ denotes all the possible permutations of the three-mode indexes. $E_{\tau}^{\mu|\nu}$ is the contangle of subsystems of $\mu$ ($\mu$ contains one mode) and $\nu$ ($\nu$ contains one or two modes), which can be defined by a proper entanglement monotone, e.g., the squared logarithmic negativity. Based on Eq. (18), the one-mode-vs-one-mode contangle $E^{r|s}_{\tau}$ and $E^{r|t}_{\tau}$ can be directly obtained by employing its definition, namely, $E_{\tau}^{\mu|\nu}\equiv[E_{\mathcal{N}}^{\mu|\nu}]^{2}$. However, when calculating the one-mode-vs-two-modes contangle $E^{r|st}_{\tau}$, one must alter the basic definition of Eq. (18) by rewriting the definition of $\eta_{0}^{-}$ as given by $\eta_{0}^{-}\equiv\min\,[\textrm{eig}|i\Omega_{3}\tilde{V}_{r|st}|]$, where $\eta_{0}^{-}$ becomes the minimum symplectic eigenvalue of the partial transpose of a $6\times 6$ CM $V$, with $\Omega_{3}=\oplus_{k=1}^{3}i\sigma_{y}$ and $\sigma_{y}$ the y-Pauli matrix. $\tilde{V}_{r|st}$ corresponds to the partial transpose of $V$, which connects to $V$ with the relation of $\tilde{V}_{r|st}=P_{r|st}VP_{r|st}$, and $P_{r|st}=\textrm{diag}(1,-1,1,1,1,1)$ is the partial transposition matrix. In terms of $E^{s|rt}_{\tau}$ and $E^{t|sr}_{\tau}$, the corresponding partial transpose matrices are given by $P_{s|rt}=\textrm{diag}(1,1,1,-1,1,1)$ and $P_{t|sr}=\textrm{diag}(1,1,1,1,1,-1)$, respectively. In addition, according to the Coffman-Kundu-Wootters monogamy inequality for quantum entanglement, we also note that the residual contangle is required to satisfy the following monogamy condition, i.e., $E^{r|st}_{\tau}-E^{r|s}_{\tau}-E^{r|t}_{\tau}\geq 0$, which means that the bipartite entanglement between the partition $r$ and the remaining two partitions $st$ is never smaller than the sum of the $r|s$ and $r|t$ bipartite entanglements in the reduced states. As such, if there are nonzero values of the minimum residual contangle, i.e., $\mathcal{R}_{\tau}^{\textrm{min}}>0$, one can verify that the full tripartite entanglement is present for the CV system.
III Nonreciprocal manipulation of quantum entanglement and EPR steering
Now we start to investigate the behaviors of quantum entanglement and EPR steering in regards to the coherent driving fields input from the CW and CCW direction in our proposed scheme. For this purpose, we evaluate the entanglement measures by numerically solving the dynamics of the steady-state CM $V$ (16) within the following parameters: $\kappa/2\pi=\textrm{1750\,Hz}$, $\omega_{m}/2\pi=\textrm{1945\,Hz}$, $\gamma_{m}/2\pi=\textrm{195\,Hz}$, $g_{1}/2\pi=\textrm{214\,Hz}$, $g_{2}/2\pi=\textrm{321\,Hz}$, $\chi/2\pi=\textrm{292\,Hz}$, $\bar{n}_{m}=\textrm{0.9}$, $\theta_{e}=\textrm{0}$, which is partially chosen from a recent experiment Pennetta et al. (2020). Note that, for simplicity and without loss of generality, we have supposed that the two mechanical modes are identical, i.e, $\omega_{m,1}=\omega_{m,2}=\omega_{m}$, $\gamma_{m,1}=\gamma_{m,2}=\gamma_{m}$, $\bar{n}_{m,1}=\bar{n}_{m,2}=\bar{n}_{m}$, while their COM coupling strength are asymmetric, i.e., $g_{2}/g_{1}=\textrm{1.5}$. This condition is convenient for producing quantum entanglement and EPR steering.
@@ III Nonreciprocal manipulation of quantum entanglement and EPR steering
Figure 3: Squeezing-induced nonreciprocal EPR steering. The quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ (a), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ (b), $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ (c), and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ (d) for opposite optical input directions are plotted as as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$. With the increase of $r_{d}$, EPR steering can be stepwise driven from no-way regime, one-way regime to two-way regime for the CCW input case but not for the CW input case, indicating the achievement of nonreciprocal EPR steering.
We first present how to achieve nonreciprocal bipartite and tripartite entanglement by leveraging a directional two-photon parametric driving field. Figure 2 shows the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ for opposite optical input directions as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$. Here $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{j}}}$ corresponds to the logarithmic negativity of the bipartition of the optical mode and the $j$th mechanical mode, while $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ corresponds to that of the bipartition of the two mechanical modes. In terms of the CW input case, as shown in the left panels of Figs. 2(a)-2(d), it is seen that in the absence of a two-photon parametric driving field, i.e., $r_{d}=r_{e}=0$, all of the logarithmic negativity $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ are null within the considered optical detuning (yellow dashed curve), indicating that there is no bipartite and tripartite entanglement in this system when only applying a coherent driving field. In contrast, for the CCW input case, it is seen that in the presence of a two-photon parametric driving field (with, e.g., squeezing strength $r_{d}=r_{e}=0.3$), the profiles of $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{1}}}$, $E_{\mathcal{N}}^{\textit{a}|\textit{m}_{\textit{2}}}$, $E_{\mathcal{N}}^{\textit{m}_{\textit{1}}|\textit{m}_{\textit{2}}}$ and $\mathcal{R}^{\textrm{min}}_{\tau}$ are all characterized by a sharp peak around the optical detuning at COM resonance $\Delta_{c}/\omega_{m}\sim\textrm{1}$. Meanwhile, as shown in the right panels of Figs. 2(a)-2(d), the bipartite and tripartite entanglement are also sensitive to the variation of the squeezing reference angle $\theta_{d}$, which all reach their maximum values at an optimal phase angle of $\theta_{d}=\pi$. This result is consistent with that of the previous study Jiao et al. , in which it reveals that the maximum COM interaction and the minimum optical noise are both obtained under the phase-matched condition with $r_{e}-r_{d}=0$ and $\theta_{e}-\theta_{d}=\pm n\pi~{}(n=1,3,5,\ldots)$, resulting in an enhanced COM entanglement. By comparing these results of the CW and CCW input cases, one can intuitively find that by applying a directional two-photon parametric driving field with a proper squeezing phase, the nonreciprocal generation and manipulation of bipartite and tripartite entanglement can be achieved, namely, such entanglement can emerge only for one specific optical input direction when reversing the coherent driving field. Physically, the mechanism behind these results is that the directional input of a two-photon parametric driving field allows one to regulate the effective COM coupling strength for the CW and CCW modes in an asymmetric way, by which one can break the time-reversal symmetry of this system. This feature of the system provides a flexible tool to tailor the parameter condition for entanglement generation based on its optical input direction. Consequently, when preparing the system under a directional entanglement generation condition, nonreciprocal bipartite and tripartite entanglement will be achieved.
@@
After having established nonreciprocal entanglement, we further show how to manipulate the directionality of EPR steering. In Fig. 3, we plot the quantum steerability $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$, $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}$ as a function of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing parameter $r_{d}$. Figures 3(a) and 3(b) show the EPR steering between the optical mode and the mechanical mode for opposite optical input directions. For the CW input case ($r_{d}=r_{e}=0$), since the optomechanical entanglement is absent, EPR steering, as a stronger quantum correlaiton, can also not be produced, where $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}$ always keep zero with the variation of the controlling parameters [see scattered dotted curve in Figs. 3(a) and 3(b)]. For the CCW input case, due to the introduction of the squeezed optical mode and the emergence of the optomechanical entanglement, one can create an asymmetric and directional EPR steering by adjusting the squeezing parameters. For instance, it is seen that although when choosing $\theta_{d}=\pi$ and some small values of $r_{d}$, such as $r_{d}=0.15$, the EPR steering still does not exist for both $\textit{a}\rightarrow\textit{m}_{\textit{1}}$ and $\textit{m}_{\textit{1}}\rightarrow\textit{a}$, i.e., $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}=\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$ [solid curve in the top panels of Figs. 3(a) and 3(b)]. However, when the squeezing strength $r_{d}$ is continued to be raised, the value of $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}$ will become nonzero and monotonically increase, resulting in an asymmetric one-way EPR steering with $\mathcal{G}^{\textit{a}\rightarrow\textit{m}_{\textit{1}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{a}}=0$. Moreover, we also show the EPR steering between the two mechanical modes for opposite optical input directions in Figs. 3(c) and 3(d). For the CW input case, there is no EPR steering for both $\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}$ and $\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}$, i.e., $\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$ [see scattered dotted curve in Figs. 3(c) and 3(d)], which is similar to the case discussed above for the optomechanical bipartition. For the CCW input case, accompanied by the occurrence of entanglement, we can get rich properties of EPR steering, where the overall state’s asymmetry is stepwise driven through the no-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}=0$), one-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}=0$ but $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$), and finally two-way regime ($\mathcal{G}^{\textit{m}_{\textit{1}}\rightarrow\textit{m}_{\textit{2}}}>0$ and $\mathcal{G}^{\textit{m}_{\textit{2}}\rightarrow\textit{m}_{\textit{1}}}>0$) with the increase of the squeezing strength $r_{d}$. From the above discussions, it is found that EPR steering can be generated and manipulated only for the CCW input case but not for the CW input case, which implies the achievement of nonreciprocal EPR steering. These results can be understood as follows: The proposed three-mode COM system has a bilinear cyclic coupling among the optical mode and the two mechanical modes. For an arbitrary bipartition $\mu$ and $\nu$, apart from their direct interaction, there is also an indirect interaction induced by their coupling to the third intermediate mode. Both direct and indirect interaction paths can generate EPR steering between $\mu$ and $\nu$, whose superposition may induce quantum interference effect that depends on the relative phase and strength of the two interaction paths. As discussed previously, by applying a directional two two-photon parametric driving field, one can manipulate the COM coupling strength in an asymmetric way. In this case, the relative strength of the two interaction paths for any bipartition of this system is influenced, based on which the manipulation of the directionality of EPR steering can be achieved by tuning the squeezing strength.

### v2

@@ II Theoretical model and system dynamics
where $(r,s,t)$ denotes all the possible permutations of the three-mode indexes. $E_{\tau}$ is the one-mode-versus-(one- or two-mode) contangle, which can be defined by a proper entanglement monotone, e.g., the squared logarithmic negativity. Based on Eq. (18), the one-mode-versus-one-mode contangle $E^{r|s}_{\tau}$ and $E^{r|t}_{\tau}$ can be directly obtained by employing its definition, namely, $E_{\tau}^{r|s}\equiv[E_{\mathcal{N}}^{r|s}]^{2}$ and $E_{\tau}^{r|t}\equiv[E_{\mathcal{N}}^{r|t}]^{2}$. However, when calculating the one-mode-versus-two-modes contangle $E^{r|st}_{\tau}$, one must alter the basic definition of Eq. (18) by rewriting the definition of $\eta_{0}^{-}$ as given by $\eta_{0}^{-}\equiv\min\,[\textrm{eig}|i\Omega_{3}\tilde{V}_{r|st}|]$, where $\eta_{0}^{-}$ becomes the minimum symplectic eigenvalue of the partial transpose of a $6\times 6$ CM $V_{r|st}$, with $\Omega_{3}=\oplus_{k=1}^{3}i\sigma_{y}$ and $\sigma_{y}$ the y-Pauli matrix. $\tilde{V}_{r|st}$ corresponds to the partial transpose of $V_{r|st}$, which connects to $V_{r|st}$ with the relation of $\tilde{V}_{r|st}=P_{r|st}VP_{r|st}$, and $P_{r|st}=\textrm{diag}(1,-1,1,1,1,1)$ is the partial transposition matrix. In terms of $E^{s|rt}_{\tau}$ and $E^{t|sr}_{\tau}$, the corresponding partial transpose matrices are given by $P_{s|rt}=\textrm{diag}(1,1,1,-1,1,1)$ and $P_{t|sr}=\textrm{diag}(1,1,1,1,1,-1)$, respectively. In addition, according to the Coffman-Kundu-Wootters monogamy inequality for quantum entanglement, we also note that the residual contangle is required to satisfy the following monogamy condition, i.e., $E^{r|st}_{\tau}-E^{r|s}_{\tau}-E^{r|t}_{\tau}\geq 0$, which means that the bipartite entanglement between the partition $r$ and the remaining two partitions $st$ is never smaller than the sum of the $r|s$ and $r|t$ bipartite entanglements in the reduced states. As such, if there are nonzero values of the minimum residual contangle, i.e., $\mathcal{R}_{\tau}^{\textrm{min}}>0$, one can verify that the full tripartite entanglement is present for the CV system.
Besides, EPR steering is an asymmetric form of quantum entanglement, where a bipartite Gaussian state characterized by CM $V_{\mu\nu}$ may be steerable from mode $\mu$ to mode $\nu$, but not vice versa. To study the quantum steerability from mode $\mu$ to mode $\nu$ or from mode $\nu$ to mode $\mu$, we adopt an intuitive and computable quantification of EPR steering [17], which are defined as
$\mathcal{S}^{\mu\rightarrow\nu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{A}_{\mu}}{4\det V_{\mu\nu}}\right],$
$\mathcal{S}^{\nu\rightarrow\mu}=$ $\max\left[0,\dfrac{1}{2}\ln\dfrac{\det\mathcal{B}_{\nu}}{4\det V_{\mu\nu}}\right].$ (21)
The quantity $\mathcal{S}$ is a monotone under Gaussian local operations and classical communication (LOCC) and the larger value of $\mathcal{S}$ implies the stronger Gaussian steerability. The steerability between mode $\mu$ and mode $\nu$ can be classified into three cases. For $\mathcal{S}^{\mu\rightarrow\nu}=\mathcal{S}^{\nu\rightarrow\mu}=0$, the state is non-steerable in any direction, which corresponds to the case of no-way steering. For $\mathcal{S}^{\mu\rightarrow\nu}>0$ and $\mathcal{S}^{\nu\rightarrow\mu}>0$, the state is steerable in both directions $\mu\rightarrow\nu$ and $\nu\rightarrow\mu$, which corresponds to the case of two-way steering. Finally, for $\mathcal{S}^{\mu\rightarrow\nu}>0$ and $\mathcal{S}^{\nu\rightarrow\mu}=0$ or $\mathcal{S}^{\nu\rightarrow\mu}>0$ and $\mathcal{S}^{\mu\rightarrow\nu}=0$, the state is only steerable in one certain direction $\mu\rightarrow\nu$ or $\nu\rightarrow\mu$, which corresponds to the case of one-way steering and reflects the asymmetric nature of entangled state.
III Phase-selective manipulation of quantum entanglement and EPR steering
First, we investigate the phase-dependent behaviors of optomechanical entanglement and the associated asymmetric EPR steering in Figs. 2 and 3. To evaluate the entanglement measures, one can numerically calculate the steady-state CM $V$ by solving the dynamical equation (16) with the following parameters: $\kappa/2\pi=4.9\,\textrm{MHz}$, $\omega_{m}/2\pi=16\,\textrm{MHz}$, $Q_{m}\equiv\omega_{m}/\gamma_{m}=10^{5}$, $G_{1}/2\pi=0.16\,\textrm{MHz}$, $G_{2}/2\pi=0.21\,\textrm{MHz}$, $\lambda/2\pi=0.32\,\textrm{MHz}$, and $\bar{n}_{m}=\textrm{100}$. Here, for ensuring the validity of the phonon-hopping interaction, we have supposed that the fundamental parameters of the two mechanical modes are identical, i.e, $\omega_{m,1}=\omega_{m,2}=\omega_{m}$, $\gamma_{m,1}=\gamma_{m,2}=\gamma_{m}$, $\bar{n}_{m,1}=\bar{n}_{m,2}=\bar{n}_{m}$. We note that these selected parameters ensure the condition required by the resolved sideband regime and mechanical ground-state cooling, i.e., $\kappa/\omega_{m}<1$, which are feasible for current experimental techniques [66] and convenient for producing optomechanical entanglement and asymmetric EPR steering. In terms of the mean thermal phonon excitation number $\bar{n}_{m}$, it corresponds to a bath temperature of $T\backsimeq 73\,\textrm{mK}$, which is achievable with the use of a dilution refrigerator. Moreover, since the effective COM coupling strength can be exponentially enhanced by the intracavity squeezing [see Fig. 1(b)], we have chosen a much smaller value for $G_{j}$ in our simulation, whose typical value can reach up to $G_{j}/2\pi\backsim 3.8\,\textrm{MHz}$ with a laser input power of $P=65\,\mu\textrm{W}$ in the near-field cavity optomechanical system  [66]. And the asymmetric COM coupling strength for the two mechanical modes can be exmperimentally operated by placing the nanomechanical oscillators at different positions along the field gradient.
In Fig. 2, we present how to achieve phase-selective generation of quantum entanglement regarding the bipartition of optical and mechanical modes. Figures 2(a) and 2(c) show the logarithmic negativity $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{2}}$ for opposite optical input directions as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$. Here $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{j}}$ corresponds to the logarithmic negativity of the bipartition of the squeezed optical mode $i$ and the mechanical mode $j$. As discussed in previous section, we have assumed the fulfillment of the phase-matched condition for the CCW input case [i.e., $\delta r_{\circlearrowleft}=\textrm{0}$ and $\delta\theta_{\circlearrowleft}=\pm n\pi(n=1,3,5...)$], whereas such condition is phase-mismatched for the CW input case [i.e., $\delta r_{\circlearrowright}\neq\textrm{0}$ and $\delta\theta_{\circlearrowright}\neq\pm n\pi~{}(n=1,3,5...)$]. Under this circumstance, it is seen that the profiles of $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{2}}$ for the CCW input case are characterized by a sharp peak around the optical detuning at COM resonance $\Delta_{c}/\omega_{m}\sim\textrm{1}$ (purple solid curve). In contrast, when reversing the coherent driving field, the values of the logarithmic negativity $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{1}}$ and $E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{2}}$ for the CW input case become smaller or even zero within the same considered optical detuning (green dashed curve). These results imply that by setting distinct squeezing phase parameters for the counterpropagating optical modes, it could lead to asymmetric response in the entanglement generation between the coherent driving field and the mechanical mode. Moreover, it is clearly seen that the response asymmetry of $E_{\mathcal{N}}^{\hat{a}_{s,i}|\hat{q}_{j}}$ is dependent on both $\delta r_{i}$ and $\delta\theta_{i}$, with stronger asymmetry attainable through appropriate optimization of the squeezing phase condition. To clearly visualize this behavior, we define the asymmetric response ratio $\mathcal{I}_{j}$ as
$\mathcal{I}_{j}=\dfrac{E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{j}}-E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowright}|\hat{q}_{j}}}{E_{\mathcal{N}}^{\hat{a}_{s,\circlearrowleft}|\hat{q}_{j}}}.$ (22)
As shown in Figs. 2(b) and 2(d), the behavior of $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are similar with the variation of the squeezing parameters, which show greater sensitivity to $\delta r_{\circlearrowright}$ than to $\delta\theta_{\circlearrowright}$ and these isolation ratios could be more than 0.9 when choosing $\delta r_{\circlearrowright}>\textrm{0.2}$ with proper values of $\delta\theta_{\circlearrowright}$. Besides, we have confirmed that the maximum values of $\mathcal{I}_{j}$ could even reach up to $1$ by properly choosing squeezing phase conditions, which means that such optomechanical entanglement can be unidirectionally produced only in one specific optical input direction when reversing the coherent driving field. Additionally, to investigate the robustness of such optomechanical entanglement against thermal noises, we further plot $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{1}}}$ and $E_{\mathcal{N}}^{\hat{\textit{a}}_{\textit{s},\textit{i}}|\hat{\textit{q}}_{\textit{2}}}$ for opposite optical input directions as functions of the mechanical environment temperature $T$ in Figs. 2(e) and 2(f). It is seen that in comparison with the CW input case, the entanglement with respect to the CCW input case is more robust against thermal noises, which can persist above a temperature of 110mK.
@@ III Phase-selective manipulation of quantum entanglement and EPR steering
Figure 4: Phase-selective generation of quantum entanglement and asymmetric EPR steering with respect to the two mechanical modes. (a),(b) The logarithmic negativity $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ with $r_{d}=0.2$ and $\theta_{d}=\pi$. The inserts show the dependence of $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and $\mathcal{R}^{\textrm{min}}_{\tau}$ on the squeezing strength difference $\delta r_{\circlearrowright}$ and the squeezing phase difference $\delta\theta_{\circlearrowright}$ at the optical detuning $\Delta_{c}/\omega_{m}=1.2$ and $1.15$, respectively. (c),(d) The measure of quantum steerability $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}$ for opposite optical input directions are plotted as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing strength $r_{d}$ with $\theta_{d}=\pi$. With the increase of $r_{d}$, EPR steering for the CCW input case can be stepwise driven from no-way regime, one-way regime to two-way regime, whereas for the CW input case, it can only be driven from no-way regime to one-way regime.
Subsequently, we further investigate how to regulate the steerability of optomechanical entanglement and achieve one-way EPR steering through tuning the squeezing phase conditions. For this purpose, we plot the measure of quantum steerability $\mathcal{S}^{\hat{a}_{s,i}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,i}}$ as functions of the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different values of $r_{d}$ in Fig. 3. It is seen that in terms of the CCW input case, for small values of $r_{d}$, both $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}}$ are null within the considered optical detuning [see red solid lines in Figs. 3(a) and 3(b)], implying that the EPR steering are not present for both directions of $\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}$ and $\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}$. When increasing the values of the intracavity squeezing strength $r_{d}$, the quantum steerability $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}$ will increase monotonically and become nonzero [see Fig. 3(a)], resulting in an asymmetric one-way EPR steering with $\mathcal{S}^{\hat{a}_{s,\circlearrowleft}\rightarrow\hat{q}_{1}}>0$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowleft}}=0$. This is because quantum steerability is an increasing function of the effective COM coupling strength that can be enhanced by the intracavity squeezing strength $r_{d}$ [see Fig. 1(c) for more details]. However, for the CW input case, since the optomechanical entanglement is absent, EPR steering, as a stronger correlated entangled state, can also not be produced, showing that $\mathcal{S}^{\hat{a}_{s,\circlearrowright}\rightarrow\hat{q}_{1}}$ and $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{a}_{s,\circlearrowright}}$ always keep zero with the variation of the controlling parameters [see green ball-solid lines in Figs. 3(a) and 3(b)].
By comparing the results of Figs. 2 and 3, it can be intuitively found that when optomechanical entanglement is prepared in an asymmetric way, the generation of the corresponding EPR steering also exhibits asymmetric characteristics, with its directionality being asymmetrically controlled via intracavity squeezing parameters. Physically, the mechanism behind these phenomena can be understood as follows: As discussed previously in Fig. 1(c), the optical input noises $N_{s,i}$ and $M_{s,i}$ can either be reduced or enhanced through adjusting the squeezing parameters $r_{e,i}$ and $\theta_{e,i}$, which provides an effective method to regulate the noise flows in CW and CCW modes in a controllable way. Here, as a specific example, we have chosen the parameter conditions where optical input noises are reduced for the CCW input case [corresponding to phase-matched condition] but amplified for the CW input case [corresponding to phase-mismatched condition]. This squeezing-phase-induced asymmetry could lead to differences in the quantum statistical properties of the CW and CCW modes. As shown in Fig. 1(d), without the presence of coherent driving, the steady-state CM $V$ displays a dominant diagonal, which is consistent with a Gaussian distribution of a vacuum state. For the CCW input case, except for the dominant diagonal elements, the steady-state CM $V$ also displays some off-diagonal elements, which are not only present in the mechanical bipartition, but also emerges in the optomechanical bipartition. For the CW input case, due to the optical input noises $N_{s,\circlearrowright}\neq 0$ and $M_{s,\circlearrowright}\neq 0$, the off-diagonal elements with respect to the optomechanical bipartition vanish in steady-state CM $V$. These results indicate that the quantum correlations between the squeezed optical mode and the mechanical modes, enabled by the down-conversion interaction through nonlinear COM coupling, survive only with small thermal photon occupancy. Therefore, by asymmetrically choosing phase conditions for the CW and CCW modes, the phase-selective generation of optomechanical entanglement and EPR steering could be achieved in this configuration.
@@
Finally, we further present how to achieve phase-selective generation of quantum entanglement and EPR steering involving two mechanical modes in Fig. 4. Given that the mechanical entanglement is more fragile to thermal noises, the parameters for numerical simulation are altered as follows: $\kappa/\textrm{2}\pi=\textrm{14.4}\,\textrm{MHz}$, $\omega_{m}/\textrm{2}\pi=\textrm{16}\,\textrm{MHz}$, $Q_{m}\equiv\omega_{m}/\gamma_{m}=\textrm{10}$, $G_{1}/\textrm{2}\pi=\textrm{1.6}\,\textrm{MHz}$, $G_{2}/\textrm{2}\pi=\textrm{2.1}\,\textrm{MHz}$, $\lambda/\textrm{2}\pi=\textrm{4.8}\,\textrm{MHz}$, and $\bar{n}_{m}=\textrm{0.9}$. Figures 4(a) and 4(b) show the logarithmic negativity $E_{\mathcal{N}}^{\hat{q}_{1}|\hat{q}_{2}}$ and the minimum residual contangle $\mathcal{R}^{\textrm{min}}_{\tau}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ for opposite optical input directions. It is seen that for the CCW input case, by fulfilling the phase-matched condition, the bipartite mechanical entanglement and the tripartite entanglement are present within a finite interval of $\Delta_{c}$ around $\Delta_{c}/\omega_{m}\simeq\textrm{1.2}$. In contrast, for the CW input case, while the bipartite mechanical entanglement and the tripartite entanglement can still persist for certain values of mismatched phase (e.g., $\delta r_{\circlearrowright}=\textrm{0.4}$ and $\delta\theta_{\circlearrowright}=\pi$), their maximum values are suppressed compared to those in the CCW input case. Moreover, from the inserts of Figs. 4(a) and 4(b), it is also seen that when setting squeezing phase difference $\delta\theta_{\circlearrowright}=\textrm{0}$, the corresponding bipartite and tripartite entanglement can completely vanish at high values of $\delta r_{\circlearrowright}$. These results are consistent with the previous discussions for optomechanical entanglement, implying the achievement of phase-selective asymmetric response in bipartite mechanical entanglement and tripartite entanglement. In Figs. 4(c) and 4(d), to explore the behaviors of EPR steering between the two mechanical modes, we further plot the quantum steerability $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}$ versus the scaled optical detuning $\Delta_{c}/\omega_{m}$ under different squeezing strength $r_{d}$ with $\theta_{d}=\pi$. For the CCW input case, it is seen that accompanied with the occurrence of highly mechanical entangled state, we can get rich properties of EPR steering, where the overall state’s asymmetry can be stepwise driven through the no-way regime ($\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}=\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}=0$), one-way regime ($\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{1}}=0$ but $\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}>0$), and finally two-way regime ($\mathcal{S}^{\hat{q}_{1}\rightarrow\hat{q}_{2}}>0$ and $\mathcal{S}^{\hat{q}_{2}\rightarrow\hat{q}_{2}}>0$) with the increase of the intracavity squeezing strength $r_{d}$. However, for the CW input case, due to the phase-mismatch induced suppression of mechanical entanglement, it is seen that the directionality of EPR steering can only be driven from no-way regime to one-way regime under the same intracavity squeezing strength. From the above discussions, it is found that the mechanical EPR steering can be generated and manipulated in an asymmetric way. These results can be understood as follows: The proposed three-mode COM system has a bilinear cyclic coupling among the optical mode and the two mechanical modes. For an arbitrary bipartition $\mu$ and $\nu$, apart from their direct interaction, there is also an indirect interaction induced by their coupling to the third intermediate mode. The superposition of direct and indirect interaction paths can give rise to a quantum interference effect, whose strength is determined by their relative phase and coupling strength. Because both quantum entanglement and EPR steering increase monotonically with the coupling strength, their magnitudes can thus be modulated via the interference effect, resulting in the generation of tripartite entanglement. In this case, by regulating the two effective COM coupling strengths $\Lambda_{j}$ through tuning the intracavity squeezing $r_{d}$, the entanglement originally arising from nonlinear COM interactions could be transferred to the two mechanical modes via the interference effect. Moreover, since we have assumed different phase conditions for the CW and CCW modes, the bipartite and tripartite entanglement will be produced in an asymmetric way when the coherent driving field is applied from opposite directions. Also, due to the asymmetric COM coupling induced for the two mechanical modes, the overall asymmetry of EPR steering can be well controlled by the squeezing phase condition.